Suppose that the New England Colonials baseball team is equally likely to win any particular game as not to win it. Suppose also that we choose a random sample of 30 Colonials games.
A. Estimate the number of games in the sample that the Colonials win by giving the mean of the relevant distribution (that is, the expectation of the relevant random variable). Do not round your response.
B. Quantify the uncertainty of your estimate by giving the standard deviation of the distribution. Round your response to at least three decimal places.

Answer by Boreal(15235)

mean is np=30*0;.5=15 games
variance is np(1-p)=15*0.5=7.5 games^2
sd is sqrt(V)=2.7386 games